OpenStudy (adi3):

for each of the following quadratic functions: (i) Find the coordinates for the vertex (ii) determine whether the vertex is max or min (iii) find the range for the functions (a) f(x) = 2x^2 +4

OpenStudy (adi3):

@SolomonZelman

OpenStudy (adi3):

@AaronAndyson

OpenStudy (adi3):

@imqwerty

OpenStudy (solomonzelman):

\(\color{black}{ \displaystyle f(x)=2x^2+4~~~~\Longrightarrow ~~~~f(x)=2(x-0)^2+4 }\) And this is in the form of \(\color{black}{ \displaystyle f(x)=2(x-h)^2+k }\) (with vertex \((h,k)\) )

OpenStudy (adi3):

ok so the vertex is 0,4

OpenStudy (solomonzelman):

yes, correct.

OpenStudy (adi3):

The vertex is min

OpenStudy (solomonzelman):

If the leading coefficient is POSITIVE, then parabola opens UP, the vertex is the MINIMUM point (or the absolute MINIMUM). If the leading coefficient is NEGATIVE, then parabola opens DOWN, the vertex is the MAXIMUM point (or the absolute MAXIMUM).

OpenStudy (solomonzelman):

Yes, so the vertex is minimum.

OpenStudy (adi3):

but you said max

OpenStudy (solomonzelman):

When the leading coefficient is negative.

OpenStudy (adi3):

ohh ok

OpenStudy (solomonzelman):

But, as I said, when leading coefficient is positive (in your case it is 2), then the vertex is minimum.

OpenStudy (solomonzelman):

And the range of the opening up parabola will start from the vertex, but has no restrictions on how large the output can be....

OpenStudy (adi3):

so it is (y:y≤4)

OpenStudy (solomonzelman):

Yes, that is exactly right.

OpenStudy (adi3):

??

OpenStudy (solomonzelman):

Or, in interval notation \({\rm y} \in [4,+\infty)\)

OpenStudy (adi3):

we did not learn that yet.

OpenStudy (solomonzelman):

All that means is that the range goes from and including 4, and ends at positive infinity (meaning that it goes up forever endlessly).

OpenStudy (solomonzelman):

don't want to throw in confusion.... in any case you are right y\(\le\)4.

OpenStudy (solomonzelman):

Any questions about what we have done?

OpenStudy (adi3):

nop, thanks a lot really apreciate, really needed it, i have a quiz tmrw.

OpenStudy (solomonzelman):

Good luck on the quiz:)

OpenStudy (adi3):

thanks mate

OpenStudy (adi3):

thanks @imqwerty for seeing my question.

imqwerty (imqwerty):

lolol XD

OpenStudy (adi3):

lol